### v3

@@ 4.6.2 Pattern recognition
The reconstruction of PLATON 3D images differs from what data processing in particle physics experiments typically requires, and is illustrated in Fig. 9b. The clustering algorithm is based on Gaussian Mixture Model (GMM) [115]: a probabilistic model that represents complex data as a combination of simpler sub-distributions, Gaussian in this case, of underlying hidden classes. The optimal number of clusters was selected using the elbow method, a heuristic that identifies the point beyond which increasing the number of clusters yields only marginal improvements in model performance. In a second step, principal component analysis (PCA) was used to merge clusters into tracks, each assigned to a different particle. This process iterated through all pairs of clusters, merging two clusters if they were both collinear and sufficiently close to each other. Once all the particle tracks of the event were obtained, the vertex of the neutrino interaction was reconstructed. Since the muon is assumed to be determined by an external detector, we use its exiting point on the scintillator surface to assign one of the tracks to the muon. Finally, the extreme of the muon track opposite to its exiting point is identified as the neutrino interaction vertex. In some cases, multiple scattering can break the muon track into different consecutive segments. To overcome this issue, the clustering algorithm is rerun with slightly relaxed conditions, potentially adding a new cluster to the muon track, until a connected but non-collinear cluster is found. If the muon does not escape the scintillator volume, the longest reconstructed track is assigned to it, and the end with the largest number of connected clusters is identified as the interaction vertex. If the two track ends exhibit an equal number of connected clusters, the point forming the smaller angle with the connected cluster is selected as the interaction vertex.
@@ 4.6.3 Selection of neutrino events
The CC sample is the most common final-state topology below 1 GeV, observed in experiments such as T2K or Hyper-Kamiokande, since it is dominated by CC quasi-elastic (CCQE) interactions. Another type of final-state topology is named 2 protons - 2 holes (2p2h) and can be induced by three main types of processes: meson exchange current, where the momentum transfer is shared between two nucleons via the exchange of a virtual meson [116]; short-range nucleon-nucleon correlations (SRC) [117], where the neutrino interacts with a nucleon that is part of a correlated nucleon-nucleon pair and both nucleons are knocked out of the nucleus; FSI, with a nucleon that knocks out a second nucleon. 2p2h events are less frequent () than CCQE interactions (), but are notoriously affected by significant systematic uncertainties in the modelling of the process [118, 119, 120]. Thus, the correct modelling of 2p2h interactions is of vital importance for the ongoing precision measurements of neutrino oscillations [121, 122, 123, 124] and, in particular, for the next-generation LBL experiments that will search for leptonic charge-parity violation [125, 3, 126].
The accurate identification of the type of each particle produced by the neutrino interaction relies on the precise measurement of the energy loss along its trajectory. For instance, the detection of the Bragg peak of particles stopping in the scintillator ensures unambiguous particle identification (PID). This feature is enhanced by the fact that most of the muons are produced with a momentum peaked around 0.5 GeV/c with a long tail above 1 GeV, while protons are expected dominantly below 1 GeV/c with a peak around 300 MeV/c. The reconstructed distribution of the energy loss along the track of protons (often stopping in the scintillator) and muons is shown in Fig. 5. To identify the particle type associated with each reconstructed track independently, a boosted decision tree (BDT) classifier was employed using the XGBoost algorithm [127]. The kinematic variables used as input features to the BDT include: the variance of the number of photons per mm, the total number of photons in the track, the maximum number of photons per mm, the minimum number of photons per mm, and the mean number of photons per mm.
@@ 4.6.4 Selection of neutrino events with a system of classical cameras
The same study has been conducted with a configuration of the PLATON module where a conventional optical system is obtained by removing the MLA. It was found that, although a system based on standard cameras exhibits inferior spatial resolution compared to an equivalent system employing plenoptic cameras (see Sec. 4.4.1), even orthogonal arrays of conventional cameras instrumented with SPAD array sensors can achieve very good performance within a PLATON module over a range of 10 cm. On the other hand, the drop in selection efficiency below 50% occurs at 215 MeV/ (corresponding to a 3.5 mm proton range) with the MLA, and at 230 MeV/ (corresponding to a 5 mm proton range) without the MLA. A similar difference can be found in the neutrino vertex resolution and the proton angular resolution: the vertex resolution drops from 0.4 mm to 0.6 mm, while the angular resolution falls from to . This is in line with the results shown in Sec. 4.4.1. Without the MLA, the efficiency and the purity are, respectively, 6% and 5% lower compared to the plenoptic configuration. It is worth noting that the PLATON module is required to cover a depth of field of 10 cm; therefore, it is plausible that eight classical cameras combined can achieve a reasonably good depth resolution.
@@ Acknowledgements
This work was supported by the Swiss National Science Foundation under grant PCEFP2_203261. This research was also partially supported by the Swiss National Science Foundation (grant 20QT21_187716 Qu3D “Quantum 3D Imaging at high speed and high resolution”). The neural network training in this work used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy User Facility, using AI4Sci@NERSC award DDR-ERCAP0034642. We would like to thank Prof. André Rubbia from ETH Zurich for useful inputs and discussions and for providing access to laboratory equipment and facilities; Prof. Vincenzo Berardi at Politecnico di Bari, Dr. Umut Kose and Johannes Wüthrich from ETH Zurich for useful discussions; Arne Erdmann from Raytrix GmbH for helping to understand the functioning of the plenoptic camera prototype assembled at Raytrix and the use of the RxLive software; Prof. Wallny at ETH Zurich for providing access to his group’s thermal chamber.

### v4

@@ 4.6.2 Pattern recognition
The reconstruction of PLATON 3D images differs from what data processing in particle physics experiments typically requires, and is illustrated in Fig. 9b. The clustering algorithm is based on Gaussian Mixture Model (GMM) [10.1111/j.2517-6161.1977.tb01600.x]: a probabilistic model that represents complex data as a combination of simpler sub-distributions, Gaussian in this case, of underlying hidden classes. The optimal number of clusters was selected using the elbow method, a heuristic that identifies the point beyond which increasing the number of clusters yields only marginal improvements in model performance. In a second step, principal component analysis (PCA) was used to merge clusters into tracks, each assigned to a different particle. This process iterated through all pairs of clusters, merging two clusters if they were both collinear and sufficiently close to each other. Once all the particle tracks of the event were obtained, the vertex of the neutrino interaction was reconstructed. Since the muon is assumed to be determined by an external detector, we use its exiting point on the scintillator surface to assign one of the tracks to the muon. Finally, the extreme of the muon track opposite to its exiting point is identified as the neutrino interaction vertex. In some cases, multiple scattering can break the muon track into different consecutive segments. To overcome this issue, the clustering algorithm is rerun with slightly relaxed conditions, potentially adding a new cluster to the muon track, until a connected but non-collinear cluster is found. If the muon does not escape the scintillator volume, the longest reconstructed track is assigned to it, and the end with the largest number of connected clusters is identified as the interaction vertex. If the two track ends exhibit an equal number of connected clusters, the point forming the smaller angle with the connected cluster is selected as the interaction vertex.
@@ 4.6.3 Selection of neutrino events
The CC sample is the most common final-state topology below 1 GeV, observed in experiments such as T2K or Hyper-Kamiokande, since it is dominated by CC quasi-elastic (CCQE) interactions. Another type of final-state topology is named 2 protons - 2 holes (2p2h) and can be induced by three main types of processes: meson exchange current, where the momentum transfer is shared between two nucleons via the exchange of a virtual meson [ALVAREZRUSO20181]; short-range nucleon-nucleon correlations (SRC) [RevModPhys.89.045002], where the neutrino interacts with a nucleon that is part of a correlated nucleon-nucleon pair and both nucleons are knocked out of the nucleus; FSI, with a nucleon that knocks out a second nucleon. 2p2h events are less frequent () than CCQE interactions (), but are notoriously affected by significant systematic uncertainties in the modelling of the process [Gran:2013kda, Nieves:2012yz, Sobczyk:2020dkn]. Thus, the correct modelling of 2p2h interactions is of vital importance for the ongoing precision measurements of neutrino oscillations [2p2h-nova, PhysRevD.105.072001, PhysRevD.96.092006, t2k-oa-2p2h-2] and, in particular, for the next-generation LBL experiments that will search for leptonic charge-parity violation [Abe:2018uyc, Abi:2020qib, DUNE:2021mtg].
The accurate identification of the type of each particle produced by the neutrino interaction relies on the precise measurement of the energy loss along its trajectory. For instance, the detection of the Bragg peak of particles stopping in the scintillator ensures unambiguous particle identification (PID). This feature is enhanced by the fact that most of the muons are produced with a momentum peaked around 0.5 GeV/c with a long tail above 1 GeV, while protons are expected dominantly below 1 GeV/c with a peak around 300 MeV/c. The reconstructed distribution of the energy loss along the track of protons (often stopping in the scintillator) and muons is shown in Fig. 5. To identify the particle type associated with each reconstructed track independently, a boosted decision tree (BDT) classifier was employed using the XGBoost algorithm [10.1145/2939672.2939785]. The kinematic variables used as input features to the BDT include: the variance of the number of photons per mm, the total number of photons in the track, the maximum number of photons per mm, the minimum number of photons per mm, and the mean number of photons per mm.
@@ 4.6.4 Selection of neutrino events with a system of classical cameras
The same study has been conducted with a configuration of the PLATON-10cm module where a conventional optical system is obtained by removing the MLA. It was found that, although a system based on standard cameras exhibits inferior spatial resolution compared to an equivalent system employing plenoptic cameras (see Sec. 4.4.1), even orthogonal arrays of conventional cameras instrumented with SPAD array sensors can achieve very good performance within a PLATON-10cm module over a range of 10 cm. On the other hand, the drop in selection efficiency below 50% occurs at 215 MeV/ (corresponding to a 3.5 mm proton range) with the MLA, and at 230 MeV/ (corresponding to a 5 mm proton range) without the MLA. A similar difference can be found in the neutrino vertex resolution and the proton angular resolution: the vertex resolution drops from 0.4 mm to 0.6 mm, while the angular resolution falls from to . This is in line with the results shown in Sec. 4.4.1. Without the MLA, the efficiency and the purity are, respectively, 6% and 5% lower compared to the plenoptic configuration. It is worth noting that the PLATON-10cm module is required to cover a depth of field of 10 cm; therefore, it is plausible that eight classical cameras combined can achieve a reasonably good depth resolution.
@@ Acknowledgements
This work was supported by the Swiss National Science Foundation under grant PCEFP2_203261. This research was also partially supported by the Swiss National Science Foundation (grant 20QT21_187716 Qu3D “Quantum 3D Imaging at high speed and high resolution”). Neural network training in this work used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy User Facility, under the AI4Sci@NERSC award DDR-ERCAP0034642. Additional training support was provided through the Swiss AI Initiative via a grant from the Swiss National Supercomputing Centre (CSCS), project ID a149, on the Alps system. We would like to thank Prof. André Rubbia from ETH Zurich for useful inputs and discussions and for providing access to laboratory equipment and facilities; Prof. Vincenzo Berardi at Politecnico di Bari, Dr. Umut Kose and Johannes Wüthrich from ETH Zurich for useful discussions; Arne Erdmann from Raytrix GmbH for helping to understand the functioning of the plenoptic camera prototype assembled at Raytrix and the use of the RxLive software; Prof. Wallny at ETH Zurich for providing access to his group’s thermal chamber.